OpenStudy (anonymous):

Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 milliliters (ml) of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (nathan917):

1) x = amount of 15% solution to add 0.30(20) + 0.15x = 0.25(x+20) 6 + 0.15x = 0.25x + 5 1 = 0.10x x = 10 add 10 mL

OpenStudy (anonymous):

Thanks :)